Nanki Corporation purchased equipment on January 1, 2016, for $631,000. In 2016 and 2017, Nanki depreciated the asset on a straight-line basis with an estimated useful life of eight years and a $7,000 residual value. In 2018, due to changes in technology, Nanki revised the useful life to a total of 4 years with no residual value. What depreciation would Nanki record for the year 2018 on this equipment? (Round your answer to the nearest dollar amount.)

Multiple Choice

  • $104,000.

  • $104,283.

  • $237,500.

  • None of these answer choices are correct.
Answer #1

Cost of equipment = $631,000

Residual value = $7,000

Useful life = 8 years

Annual depreciation = ( Cost of equipment - Residual value ) / Useful life

= ( 631,000-7,000)/8

= $78,000

Accumulated depreciation on January 1, 2018 = 78,000 x 2

= $156,000

Book value of assets on January 1, 2018 = Cost of equipment - Accumulated depreciation on January 1, 2018

= 631,000-156,000

= $475,000

Revised useful life = 4 years

Remaining useful life = Revised useful life - years already used

= 4-2

= 2

Revised annual depreciation = Book value of assets on January 1, 2018 / Remaining useful life

= 475,000/2

= $237,500

Third option is correct option.

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks